Form the PDE by eliminating the arbitrary function from 𝑓 ( π‘₯βˆ’π‘¦ π‘¦βˆ’π‘§ , π‘₯𝑦 + 𝑦𝑧 + 𝑧π‘₯) = 0.


1
Expert's answer
2021-06-15T13:46:47-0400

"𝑓 ( \frac{x-y}{y-z}, 𝑥𝑦 + 𝑦𝑧 + 𝑧𝑥) = 0\\
let\space u=\frac{x-y}{y-z}, v=𝑥𝑦 + 𝑦𝑧 + 𝑧𝑥\\
Therefore, f(u,v)=0\\
u_x=\frac{1}{y-z}, \\u_y=\frac{-y+z-xq-yq}{(y-z)^2},\\ v_x=x+p,\\v_y=x+q\\
Now,\\
f_x=f_u.u_x+f_v.v_x=0\\
0=f_u.(\frac{1}{y-z})
+f_v.(x+p)\\
f_u.(\frac{1}{y-z})=-f_v.(x+p)
-----(1)\\
f_y=f_u.u_y+f_v.v_y=0\\
0=f_u.(\frac{-y+z-xq-yq}{(y-z)^2})
+f_v.(x+q)\\
f_u.(\frac{-y+z-xq-yq}{(y-z)^2})=-f_v.(x+q)--(2)
\\
\text{Divide equation 1 by 2, we get,}\\
\frac{\frac{1}{y-z}}{\frac{-y+z-xq-yq}{(y-z)^2}}=\frac{x+p}{x+q}\\
\implies \frac{y-z}{-y+z-xq-yq
}=\frac{x+p}{x+q}\\
\text{This is the required solution.}"